OpenStudy (anonymous):

Mrs. Collins is at the table with you and states that the fourth-degree graphs she has seen have 4 real zeros. She asks you if it is possible to create a fourth-degree polynomial with only 2 real zeros. Demonstrate how to do this and explain your steps.

OpenStudy (anonymous):

I know how to solve the equations and all, but I really dont know the first thing to creating one. Please help x.x

OpenStudy (tkhunny):

Zero of 'a' means a factor of (x-a). Four real zeros, a, b, c, d means factors of (x-a)(x-b)(x-c)(x-d). What happens if a = b and c = d?

OpenStudy (anonymous):

wow that is really confusing and I have no idea

OpenStudy (tkhunny):

What's confusing about this? Zero of 'a' means a factor of (x-a). Examples: Zero of 2 means a factor of (x-2). Zero of 7 means a factor of (x-7). Zero of -3 means a factor of (x-(-3)) = (x+3). Zero 0f zero (0) means a factor of (x-0) = x. You can do that.

OpenStudy (anonymous):

Oh okay, I just got mixed up because it was in the form of (x-a). Im used to it being x = a. I got it, thank you!